Determine the local minimum and local maximum value of the function f defined by f(x)=3-5x^3+5x^4-x^5


1
Expert's answer
2021-05-07T11:44:20-0400

Solution:

"f(x)=3-5x^3+5x^4-x^5"

"\Rightarrow f'(x)=-15x^2+20x^3-5x^4"

Put "f'(x)=0"

"\Rightarrow -15x^2+20x^3-5x^4=0
\\\Rightarrow -5x^2(3-4x+x^2)=0"

"\\\Rightarrow -5x^2=0,(3-4x+x^2)=0"

"\\\Rightarrow x=0,(3-3x-x+x^2)=0
\\\Rightarrow(3-x)(1-x)=0
\\\Rightarrow x=1,x=3,x=0"

"\mathrm{If\:}f\:'\left(x\right)>0\mathrm{\:to\:the\:left\:of\:}x=c\mathrm{\:and\:}f\:'\left(x\right)<0\mathrm{\:to\:the\:right\:of\:}x=c\mathrm{\:then\:}x=c\mathrm{\:is\:a\:local\:maximum.}"

We have, "f'(x)>0" when "1<x<3" and "f'(x)<0" when "x<0, 0<x<1,x>3", so "x=3" is a point of local maximum and "x=1" is a point of local minimum.

So, put "x=3" in "f(x)".

"f(3)=3-5(3)^3+5(3)^4-(3)^5=30"

And put "x=1" in "f(x)".

"f(1)=3-5(1)^3+5(1)^4-(1)^5=2"

Hence, local maximum is 30 at x=3 and minimum is 2 at x=1.